• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Z.-E. Meziani

    Log entry time 08:15:01 on November13,2008

    Entry number 248530

    keyword=Shift Summary

    12:00 am start of shift
    Production in run 3308

    time Run # # events target comment
    23:22-00:10 3308 998837 3He [pol-T in plane] good
    00:11-00:46 3309 998433 3He [pol-T in plane] good
    00:47-01:19 3310 1000672 3He [pol-T in plane] good
    request to lower the current to 1uA
    01:24-02:35 3311 Cherenkov voltages also lowered after looing at the scope and seeing clear saturation for the most of the signals.
    For run 3311 current was 1uA

    02:38 3312, I= 8 uA Coda crashed and rebooted
    ROC 9 frozen
    time Run # # events target comment
    03:04-03:32 3313 1000204 3He [pol-T in plane] good
    03:33-04:05 3314 999605 3He [pol-T in plane] Cer H voltage
    04:16-04:51 3315 999541 3He [pol-T in plane] Cer L voltage
    04:52-05:19 3316 1000983 3He [pol-T in plane] Cer H voltage
    05:20-05:51 3317 1000075 3He [pol-T in plane]
    05:52-05:55 3318 98842 3He [pol-T in plane] Stopped to include the RICH detector
    06:03-06:08 3319 141061 3He [pol-T in plane] stopped early to change prescaler (deadtime was 22%)
    06:19-06:49 3320 999076 3He [pol-T in plane]
    06:52-06:54 3321 bad run
    06:55-07:37 3322 999160 3He[pol-T in plane] good
    07:38- 3323 Cosmics

    End of shift 8:00 am.